ABC Company purchased a mine in 2017. It was estimated that the mine contained 250,000 tons of ore and that the mine would be worthless after all of the ore was extracted. The company extracted 25,500 tons of ore in 2017 and 30,000 tons of ore in 2018. In 2017, ABC Company sold 22,000 tons of ore. In 2018, the company sold 15,000 tons of ore. Accumulated depletion at December 31, 2018 was $999,000. What was original cost of the mine?
Answer #1

Total tons of ore extracted during 2017 & 2018 = 25500+30000 = 55,500

Accumulated depletion at December 31, 2018 = $999,000

Depletion rate per ton = Accumulated depletion at December 31, 2018 / Total tons of ore extracted during 2017 & 2018

= $999,000 / 55,500

= $18 per ton

Original cost of the mine = Total ore contained in mine x Depletion rate per ton

= 250,000*18

= $4500000